Factor by grouping x^2-y^2-ax-ay
Found 3 solutions by mananth, ikleyn, n2:
Answer by mananth(16949)

x^2-y^2-ax-ay
(x+y)(x-y) - a(x+y)
(x+y){ x-y -ax+ay)
(x+y){ x-ax-y+ay}
(x+y){x(1-a)-y(1-a)}
(x+y){(1-a)(x-y)}
(x+y)(x-y)(1+a)

Answer by ikleyn(53875)

.
Factor by grouping x^2-y^2-ax-ay
~~~~~~~~~~~~~~~~~~~~~~~~


        In the post by @mananth, factoring is made incorrectly, giving wrong answer.
        I came to bring a correct solution.


   x^2-y^2-ax-ay = 

= (x+y)*(x-y) - a(x+y)

= (x+y)*{x-y-a)


This is the final ANSWER:  x^2-y^2-ax-ay = (x+y)*(x-y-a).

Solved correctly.



Answer by n2(88)

.
Factor by grouping x^2-y^2-ax-ay
~~~~~~~~~~~~~~~~~~~~~~~~


   x^2-y^2-ax-ay = 

= (x+y)(x-y) - a(x+y)

= (x+y){ x-y -a)


This is the final ANSWER:  x^2-y^2-ax-ay = (x+y)*(x-y-a).

Solved.